# Totally elliptic surface group representations

### v1

@@ 2.3. Genus zero
In order to finish the proof of Theorem A, we shall prove that a reduced totally elliptic genus-0 surface group representation $\rho\colon\pi_{1}\Sigma\to\operatorname{PSL}_{2}\operatorname{\mathbb{R}}$ is either orthogonal or a DT representation. Clearly, if $\Sigma$ has at most two punctures, then $\rho$ is orthogonal. Let’s now assume that $\Sigma$ is a sphere with a set $\mathcal{P}$ of $n\geq 3$ punctures. Recall that the simple closed curves on $\Sigma$ that are homotopic to a puncture are called peripheral. By definition, a reduced totally elliptic representation maps every peripheral curve on $\Sigma$ to a regular elliptic element. In other words, the conjugacy class $[\rho]$ of a reduced totally elliptic representation $\rho\colon\pi_{1}\Sigma\to\operatorname{PSL}_{2}\operatorname{\mathbb{R}}$ belongs to an $\alpha$-relative character variety
$\operatorname{Rep}_{\alpha}(\Sigma,\operatorname{PSL}_{2}\operatorname{\mathbb{R}})$
which consists of all conjugacy classes of representations mapping peripheral curves around every puncture $p\in\mathcal{P}$ to regular elliptic elements in $\operatorname{PSL}_{2}\operatorname{\mathbb{R}}$ of fixed rotation angle $\alpha_{p}\in(0,2\pi)$. The numbers $\alpha_{p}$ are the components of the angle vector $\alpha\in(0,2\pi)^{\mathcal{P}}$. We’ll write
@@
Unlike in the positive genus case, there are Zariski dense totally elliptic representations $\pi_{1}\Sigma\to\operatorname{PSL}_{2}\operatorname{\mathbb{R}}$ when $\Sigma$ is an $n$-punctured sphere. The first examples were discovered via computer simulations by Benedetto–Goldman in the case $n=4$. Their construction was later generalized by Deroin–Tholozan for spheres with an arbitrary number of punctures.
@@
If $|\alpha|<2\pi$ or $|\alpha|>2\pi(n-1)$, then $\operatorname{Rep}_{\alpha}(\Sigma,\operatorname{PSL}_{2}\operatorname{\mathbb{R}})$ contains a compact connected component isomorphic to $\mathbb{CP}^{n-3}$ and entirely made of conjugacy classes of Zariski dense totally elliptic representations that don’t contain any simple closed curve in their kernel.
@@
According to Theorem 2.9, DT representations are more than just totally elliptic in the sense of Definition 1.1; they are “regularly totally elliptic” by which we mean that every simple closed curve is mapped to a regular elliptic element of $\operatorname{PSL}_{2}\operatorname{\mathbb{R}}$.
DT representations can be characterized by their Toledo number in the sense of Burger–Iozzi–Wienhard [BIW10]. Namely, according to Deroin–Tholozan’s original definition, a conjugacy class of representations in $\operatorname{Rep}_{\alpha}(\Sigma,\operatorname{PSL}_{2}\operatorname{\mathbb{R}})$ belongs to a DT component if and only if its Toledo number is
@@
It’s worth noticing that the Toledo number of a DT representation is always contained in the interval $(-1,1)\setminus\{0\}$ and can be arbitrarily close to 00, highlighting another strong contrast with Fuchsian representations.
The topology of relative $\operatorname{PSL}_{2}\operatorname{\mathbb{R}}$ character varieties has been studied by Mondello. Among other things, he characterized all their compact components, proving that the only compact connected components that aren’t isolated points are DT components.
@@
The relative character variety $\operatorname{Rep}_{\alpha}(\Sigma,\operatorname{PSL}_{2}\operatorname{\mathbb{R}})$ contains at most one compact connected component and it does precisely in the following cases:
@@
  • $|\alpha|\in(0,2\pi)\cup(2\pi(n-1),2\pi n)$ and the compact component is the DT component from Theorem 2.9.
In the context of punctured spheres, Theorem A can now be reformulated as follows. If an element $[\rho]\in\operatorname{Rep}_{\alpha}(\Sigma,\operatorname{PSL}_{2}\operatorname{\mathbb{R}})$ is the conjugacy class of a reduced totally elliptic representation, then $\alpha$ is one of the angle vectors from Corollary 2.11 and $[\rho]$ belongs to the unique compact connected component of $\operatorname{Rep}_{\alpha}(\Sigma,\operatorname{PSL}_{2}\operatorname{\mathbb{R}})$. Actually, Deroin–Tholozan already observed that totally elliptic representations belong to a bounded region of $\operatorname{Rep}_{\alpha}(\Sigma,\operatorname{PSL}_{2}\operatorname{\mathbb{R}})$.
@@
The closure of the subset of all conjugacy classes of totally elliptic representations inside $\operatorname{Rep}_{\alpha}(\Sigma,\operatorname{PSL}_{2}\operatorname{\mathbb{R}})$ is compact.
Deroin–Tholozan’s argument to prove Proposition 2.12 uses that all the representations in a sequence of totally elliptic representations are dominated by the same Fuchsian representation and therefore correspond to a sequence of Lipschitz equivariant maps $\operatorname{\mathbb{H}}\to\operatorname{\mathbb{H}}$ by a result of Guéritaud–Kassel [GK17]. The conclusion then follows from the Arzelà–Ascoli theorem (see [DT19] for more details).
Even though totally elliptic representations are contained in a bounded region, it could nevertheless be possible a priori to have a compact subset of a non-compact component of $\operatorname{Rep}_{\alpha}(\Sigma,\operatorname{PSL}_{2}\operatorname{\mathbb{R}})$ made of totally elliptic representations. To finish the proof of Theorem A, we shall prove that no such compact subsets exist. We’ll do so by going over all possible values of $n$ inductively (recall that $n$ is the number of punctures of $\Sigma$).
@@ 2.3.1. The case $n=3$
A $3$-punctured sphere $\Sigma$ only contains three non-trivial simple closed curves: the three peripheral curves. So any point in $\operatorname{Rep}_{\alpha}(\Sigma,\operatorname{PSL}_{2}\operatorname{\mathbb{R}})$ is automatically the conjugacy class of a totally elliptic representation. It turns out that when $n=3$, the relative character variety $\operatorname{Rep}_{\alpha}(\Sigma,\operatorname{PSL}_{2}\operatorname{\mathbb{R}})$ is either empty or a singleton. More precisely, let’s pick a geometric presentation of $\pi_{1}\Sigma$ with generators $c_{1},c_{2},c_{3}$ satisfying $c_{1}c_{2}c_{3}=1$. Now, if $\rho$ is a reduced totally elliptic representation, then the unique fixed points of $\rho(c_{1})$, $\rho(c_{2})$, and $\rho(c_{3})$ form a triangle in $\operatorname{\mathbb{H}}$ with vertices $(C_{1},C_{2},C_{3})$. Denote the rotation angles of $\rho(c_{1})$, $\rho(c_{2})$, and $\rho(c_{3})$ by $\alpha=(\alpha_{1},\alpha_{2},\alpha_{3})\in(0,2\pi)^{3}$. When we study the location of the points $(C_{1},C_{2},C_{3})$, we observe that the triangle they form can only be in three possible configurations, as explained in [Mar24, Section 3.1 and Table 1].
@@
If $|\alpha|\in\{2\pi,4\pi\}$, then $\rho$ is orthogonal. When $|\alpha|<2\pi$ or when $|\alpha|>4\pi$, then $\rho$ is a DT representation. (Note that when $|\alpha|\in(2\pi,4\pi)$, then $\operatorname{Rep}_{\alpha}(\Sigma,\operatorname{PSL}_{2}\operatorname{\mathbb{R}})$ is empty.) So, Theorem A holds for 3-punctured spheres.
@@ 2.3.2. The case $n=4$
Throughout this section, $\Sigma$ will denote a $4$-punctured sphere. We start by considering the case of representations with a non-peripheral simple closed curve in their kernel.
@@
We work with a geometric presentation of $\pi_{1}\Sigma$ with generators $c_{1},c_{2},c_{3},c_{4}$ satisfying $c_{1}c_{2}c_{3}c_{4}=1$. Let’s assume that $c_{1}c_{2}$ is a fundamental group element that represents the non-peripheral simple closed curve in the kernel of $\rho$. This means that there exist two elements $X,Y$ of $\operatorname{PSL}_{2}\operatorname{\mathbb{R}}$ such that $\rho(c_{1})=X^{-1}$, $\rho(c_{2})=X$, $\rho(c_{3})=Y$, and $\rho(c_{4})=Y^{-1}$. Now, observe that the fundamental group element $c_{2}c_{3}c_{1}c_{3}^{-1}$ represents a simple closed curve on $\Sigma$ (it’s the image of the simple closed curve represented by $c_{1}c_{2}$ under the Dehn twist along $c_{2}c_{3}$). Since $\rho$ is totally elliptic, we conclude that $\rho(c_{2}c_{3}c_{1}c_{3}^{-1})=[X,Y]$ is elliptic. By Lemma 2.1, this is possible only if $[X,Y]=1$, proving that $\rho$ is orthogonal. ∎
It was already observed in [Mar22, Remark 2.8] that if $|\alpha|<2\pi$ or if $|\alpha|>6\pi$, then every totally elliptic representation in $\operatorname{Rep}_{\alpha}(\Sigma,\operatorname{PSL}_{2}\operatorname{\mathbb{R}})$ is a DT representation. It turns out that a stronger statement is true: a representation that maps a single non-peripheral simple closed curve to a regular elliptic element is necessarily a DT representation. The proof of this statement uses the characterization of DT representations in terms of Toledo number (given by (2.3)) and the additive properties of the Toledo number (see [Mar22] for more details). When $|\alpha|\in[2\pi,6\pi]$, however, there exist representations outside of the compact component of $\operatorname{Rep}_{\alpha}(\Sigma,\operatorname{PSL}_{2}\operatorname{\mathbb{R}})$ that map some non-peripheral simple closed curves to elliptic elements, but are not totally elliptic (an explicit example is provided by Palesi in [Pal14, Section 5.4]).
We now give a more general argument that applies for every value of $|\alpha|$ and proves that totally elliptic representations always belong to a compact component of $\operatorname{Rep}_{\alpha}(\Sigma,\operatorname{PSL}_{2}\operatorname{\mathbb{R}})$. Let $\rho\colon\pi_{1}\Sigma\to\operatorname{PSL}_{2}\operatorname{\mathbb{R}}$ be a totally elliptic representation and consider the pure mapping class group orbit of $[\rho]$ inside $\operatorname{Rep}_{\alpha}(\Sigma,\operatorname{PSL}_{2}\operatorname{\mathbb{R}})$ (we refer the reader to [Mar22] for a detailed description of this action). The orbit of $[\rho]$ is a bounded subset of $\operatorname{Rep}_{\alpha}(\Sigma,\operatorname{PSL}_{2}\operatorname{\mathbb{R}})$ by Proposition 2.12. Bounded orbits for 4-punctured spheres have been studied by Cantat–Loray in the more general context of $\mathrm{SL}_{2}\operatorname{\mathbb{C}}$ representations. Their work has the following implication.
@@
Let $\Sigma$ be a 4-punctured sphere. The relative character variety $\operatorname{Rep}_{\alpha}(\Sigma,\operatorname{PSL}_{2}\operatorname{\mathbb{R}})$ contains an infinite bounded mapping class group orbit if and only if $|\alpha|<2\pi$ or $|\alpha|>6\pi$. Moreover, each infinite bounded orbit is dense in the DT component of $\operatorname{Rep}_{\alpha}(\Sigma,\operatorname{PSL}_{2}\operatorname{\mathbb{R}})$.
This means that if the mapping class group orbit of $[\rho]$ is infinite, then $|\alpha|<2\pi$ or $|\alpha|>6\pi$, and $\rho$ is a DT representation. If instead the orbit of $[\rho]$ is finite, then we can consult Lisovyy–Tykhyy’s classification of finite mapping class group orbits for 4-punctured spheres from [LT14] and observe that $\rho$ is either a representation into $\operatorname{PSO}(2)$ (which corresponds to the isolated points of $\operatorname{Rep}_{\alpha}(\Sigma,\operatorname{PSL}_{2}\operatorname{\mathbb{R}})$ when $|\alpha|\in\{2\pi,4\pi,6\pi\}$) or a DT representation. Altogether, this proves Theorem A for $4$-punctured spheres.

### v2

@@ 2.3. Genus zero
First, note that if $\Sigma$ has at most two punctures, then every totally elliptic surface group representation $\rho\colon\pi_{1}\Sigma\to\operatorname{PSL}_{2}\operatorname{\mathbb{R}}$ is orthogonal. So we will assume that $\Sigma$ is a sphere with a set $\mathcal{P}$ of $n\geq 3$ punctures.
For every angle vector $\alpha\in(0,2\pi)^{\mathcal{P}}$, the subspace of the character variety $\operatorname{Rep}(\Sigma,\operatorname{PSL}_{2}\operatorname{\mathbb{R}})$ consisting of all conjugacy classes of representations mapping every peripheral curve around a puncture $p\in\mathcal{P}$ to regular elliptic elements of fixed rotation angle $\alpha_{p}\in(0,2\pi)$ is called $\alpha$-relative character variety and is denoted by
$\operatorname{Rep}_{\alpha}(\Sigma,\operatorname{PSL}_{2}\operatorname{\mathbb{R}})\subset\operatorname{Rep}(\Sigma,\operatorname{PSL}_{2}\operatorname{\mathbb{R}}).$
The numbers $\alpha_{p}$ are called peripheral angles and their sum is written
@@
By definition, a reduced totally elliptic representation $\rho\colon\pi_{1}\Sigma\to\operatorname{PSL}_{2}\operatorname{\mathbb{R}}$ maps every peripheral curve on $\Sigma$ to a regular elliptic element of $\operatorname{PSL}_{2}\operatorname{\mathbb{R}}$. The images under $\rho$ of all peripheral curves around a puncture $p\in\mathcal{P}$ are conjugate and thus have the same rotation angle $\alpha_{p}\in(0,2\pi)$. In other words, the conjugacy class $[\rho]$ of $\rho$ belongs to the $\alpha$-relative character variety where $\alpha$ is the vector of peripheral angles $\alpha_{p}$.
In contrast to the positive genus case (see Propositions 2.5 and 2.8), there exist totally elliptic genus-0 surface group representations with Zariski dense image in $\operatorname{PSL}_{2}\operatorname{\mathbb{R}}$. The first examples were discovered via computer simulations by Benedetto–Goldman in the case $n=4$. Their construction was later generalized by Deroin–Tholozan for spheres with an arbitrary number of punctures, leading to the following combined statement.
@@
If $|\alpha|<2\pi$ or $|\alpha|>2\pi(n-1)$, then $\operatorname{Rep}_{\alpha}(\Sigma,\operatorname{PSL}_{2}\operatorname{\mathbb{R}})$ contains a compact connected component isomorphic to $\mathbb{CP}^{n-3}$ and entirely made of conjugacy classes of totally elliptic representations with Zariski dense image. More precisely, these representations map every simple closed curves on $\Sigma$ to a regular elliptic element.
@@
According to the last statement of Theorem 2.9, DT representations are more than just totally elliptic in the sense of Definition 1.1; they are “regularly totally elliptic” by which we mean that every simple closed curve is mapped to a regular elliptic element of $\operatorname{PSL}_{2}\operatorname{\mathbb{R}}$.
DT representations can be characterized by their Toledo number in the sense of Burger–Iozzi–Wienhard [BIW10]. Namely, according to Deroin–Tholozan’s original definition, a point in $\operatorname{Rep}_{\alpha}(\Sigma,\operatorname{PSL}_{2}\operatorname{\mathbb{R}})$ belongs to a DT component if and only if its Toledo number is
@@
It is worth noticing that the Toledo number of a DT representation is always contained in $(-1,1)\setminus\{0\}$ and can be arbitrarily close to 00 when $|\alpha|$ approaches $2\pi$ or $2\pi(n-1)$. Since the absolute value of the Toledo number on $\operatorname{Rep}(\Sigma,\operatorname{PSL}_{2}\operatorname{\mathbb{R}})$ can be as high as $-\chi(\Sigma)=n-2$ according to [BIW10], DT representations exhibit nearly minimal Toledo number among all points of $\operatorname{Rep}(\Sigma,\operatorname{PSL}_{2}\operatorname{\mathbb{R}})$. This highlights another strong contrast with Fuchsian representations which have maximal Toledo (or Euler) number.
The topology of relative $\operatorname{PSL}_{2}\operatorname{\mathbb{R}}$ character varieties has been studied by Mondello. Among other things, he characterized all their compact components, proving that the only compact connected components that are not isolated points are DT components.
@@
The $\alpha$-relative character variety $\operatorname{Rep}_{\alpha}(\Sigma,\operatorname{PSL}_{2}\operatorname{\mathbb{R}})$ contains at most one compact connected component and it does precisely in the following cases:
@@
  • $|\alpha|\in(0,2\pi)\cup(2\pi(n-1),2\pi n)$ and the compact component is one of the DT components from Theorem 2.9.
With Corollary 2.11 in mind and in order to conclude the proof of Theorem A, we shall show that if an element $[\rho]\in\operatorname{Rep}_{\alpha}(\Sigma,\operatorname{PSL}_{2}\operatorname{\mathbb{R}})$ is the conjugacy class of a reduced totally elliptic representation, then $\alpha$ is one of the angle vectors from Corollary 2.11 and $[\rho]$ belongs to the unique compact connected component of $\operatorname{Rep}_{\alpha}(\Sigma,\operatorname{PSL}_{2}\operatorname{\mathbb{R}})$. We will use a result of Deroin–Tholozan who observed that conjugacy classes of totally elliptic representations remain in a bounded region of the character variety.
@@
The subset of all conjugacy classes of totally elliptic representations inside the character variety $\operatorname{Rep}(\Sigma,\operatorname{PSL}_{2}\operatorname{\mathbb{R}})$ is bounded (i.e. contained in a compact subset).
Deroin–Tholozan’s argument to prove Proposition 2.12 goes as follows. First, consider the closure of the subset of totally elliptic representations inside $\operatorname{Hom}(\pi_{1}\Sigma,\operatorname{PSL}_{2}\operatorname{\mathbb{R}})$ and call it $\overline{\mathcal{TE}}$. The set $\overline{\mathcal{TE}}$ is the collection of all representations in $\operatorname{Hom}(\pi_{1}\Sigma,\operatorname{PSL}_{2}\operatorname{\mathbb{R}})$ that map every simple closed curve on $\Sigma$ to either an elliptic or a parabolic element of $\operatorname{PSL}_{2}\operatorname{\mathbb{R}}$. All the representations in $\overline{\mathcal{TE}}$ are dominated (in the sense of [GKW15, DT16]) by every Fuchsian representation in $\operatorname{Hom}(\pi_{1}\Sigma,\operatorname{PSL}_{2}\operatorname{\mathbb{R}})$ by a result of Guéritaud–Kassel [GK17]. So, if we fix an arbitrary Fuchsian representation $j\colon\pi_{1}\Sigma\to\operatorname{PSL}_{2}\operatorname{\mathbb{R}}$, it will dominate every representation in a sequence $(\rho_{k})\subset\overline{\mathcal{TE}}$. Domination means that to every $\rho_{k}$ corresponds a $(j,\rho_{k})$-equivariant 1-Lipschitz map $\operatorname{\mathbb{H}}\to\operatorname{\mathbb{H}}$. The existence of a converging subseqeunce for $(\rho_{k})$ now follows from the Arzelà–Ascoli theorem (see [DT19] for more details). This argument shows that $\overline{\mathcal{TE}}\subset\operatorname{Hom}(\pi_{1}\Sigma,\operatorname{PSL}_{2}\operatorname{\mathbb{R}})$ is sequentially compact, hence compact because $\operatorname{Hom}(\pi_{1}\Sigma,\operatorname{PSL}_{2}\operatorname{\mathbb{R}})$ is a metrizable space. Now, if $p$ denotes the quotient map from $\operatorname{Hom}(\pi_{1}\Sigma,\operatorname{PSL}_{2}\operatorname{\mathbb{R}})$ to the non-Hausdorff quotient $\operatorname{Rep}(\Sigma,\operatorname{PSL}_{2}\operatorname{\mathbb{R}})$, then the subset of all conjugacy classes of totally elliptic representations in $\operatorname{Rep}(\Sigma,\operatorname{PSL}_{2}\operatorname{\mathbb{R}})$ is contained in the compact set $p(\overline{\mathcal{TE}})$.
Even though totally elliptic representations remain in a bounded region by Proposition 2.12, it could nevertheless be possible that some non-compact components of $\operatorname{Rep}_{\alpha}(\Sigma,\operatorname{PSL}_{2}\operatorname{\mathbb{R}})$ contain conjugacy classes of totally elliptic representations. To conclude the proof of Theorem A, we shall prove that this is not the case. We will do so using an induction on the number $n$ of punctures of $\Sigma$.
@@ 2.3.1. The case $n=3$
A $3$-punctured sphere $\Sigma$ only contains three non-trivial simple closed curves: its three peripheral curves. Any point in an $\alpha$-relative character variety $\operatorname{Rep}_{\alpha}(\Sigma,\operatorname{PSL}_{2}\operatorname{\mathbb{R}})$ is thus automatically the conjugacy class of a totally elliptic representation. It turns out that when $n=3$, an $\alpha$-relative character variety $\operatorname{Rep}_{\alpha}(\Sigma,\operatorname{PSL}_{2}\operatorname{\mathbb{R}})$ is either empty or a singleton. More precisely, let us pick a geometric presentation of $\pi_{1}\Sigma$ with generators $c_{1},c_{2},c_{3}$ satisfying $c_{1}c_{2}c_{3}=1$, similarly as on Figure 1. Now, if $\rho$ is a reduced totally elliptic representation, then the unique fixed points of $\rho(c_{1})$, $\rho(c_{2})$, $\rho(c_{3})$ form a triangle in the hyperbolic plane with vertices $(C_{1},C_{2},C_{3})$. Denote the rotation angles of $\rho(c_{1})$, $\rho(c_{2})$, $\rho(c_{3})$ by $\alpha=(\alpha_{1},\alpha_{2},\alpha_{3})\in(0,2\pi)^{3}$. When we study the location of the points $(C_{1},C_{2},C_{3})$, we observe that the triangle they form can only be in three possible configurations, as explained in [Mar24, Section 3.1 and Table 1].
@@
If $|\alpha|\in\{2\pi,4\pi\}$, then $\rho$ is orthogonal. When $|\alpha|<2\pi$ or when $|\alpha|>4\pi$, then $\rho$ is a DT representation. Finally, if $|\alpha|\in(2\pi,4\pi)$, then $\operatorname{Rep}_{\alpha}(\Sigma,\operatorname{PSL}_{2}\operatorname{\mathbb{R}})$ is empty. Altogether, this proves Theorem A for 3-punctured spheres.
@@ 2.3.2. The case $n=4$
Throughout this section, $\Sigma$ will denote a sphere with four punctures. We start by considering the case of representations with a non-peripheral simple closed curve in their kernel.
@@
We work with a geometric presentation of $\pi_{1}\Sigma$ with generators $c_{1},c_{2},c_{3},c_{4}$ satisfying $c_{1}c_{2}c_{3}c_{4}=1$. Up to cyclically permuting the generators, we may assume that $c_{1}c_{2}$ is a fundamental group element that represents the non-peripheral simple closed curve in the kernel of $\rho$. This means that there exist two elements $X,Y$ of $\operatorname{PSL}_{2}\operatorname{\mathbb{R}}$ such that $\rho(c_{1})=X^{-1}$, $\rho(c_{2})=X$, $\rho(c_{3})=Y$, and $\rho(c_{4})=Y^{-1}$.
We now claim that the fundamental group element $c_{1}c_{3}^{-1}c_{2}c_{3}$ represents a simple closed curve on $\Sigma$. This can be seen directly on Figure 1 by applying the concatenation conventions of Remark 2.3. Alternatively, apply a Dehn twist along the simple closed curve represented by $c_{2}c_{3}$ to the simple closed curve represented by $c_{1}c_{2}$. The resulting closed curve is simple and represented by $c_{1}c_{3}^{-1}c_{2}c_{3}$.
Finally, using that $\rho$ is totally elliptic and that $c_{1}c_{3}^{-1}c_{2}c_{3}$ represents a simple closed curve, we conclude that $\rho(c_{1}c_{3}^{-1}c_{2}c_{3})=[X^{-1},Y^{-1}]$ is elliptic. By Lemma 2.1, this is possible only if $[X^{-1},Y^{-1}]=1$, proving that $\rho$ is orthogonal. ∎
Having proved Lemma 2.13, we will now consider reduced totally elliptic representations $\rho\colon\pi_{1}\Sigma\to\operatorname{PSL}_{2}\operatorname{\mathbb{R}}$ with no simple closed curve in their kernel. In the terminology introduced after Definition 2.10, they are regularly totally elliptic. In particular, each conjugacy class $[\rho]$ belongs to an $\alpha$-relative character variety $\operatorname{Rep}_{\alpha}(\Sigma,\operatorname{PSL}_{2}\operatorname{\mathbb{R}})$ for some angle vector $\alpha\in(0,2\pi)^{4}$. We want to prove that every such $\rho$ is either a DT or an orthogonal representation.
Let us start with the case $|\alpha|<2\pi$ or $|\alpha|>6\pi$ which was already treated in [Mar22]. It turns that mapping a single non-peripheral closed curve to an elliptic element is enough to be a DT representation in that case.
Lemma 2.14 ([Mar22, Remark 2.8]).
Let $\alpha\in(0,2\pi)^{4}$ be an angle vector such that $|\alpha|<2\pi$ or $|\alpha|>6\pi$ and let $[\rho]\in\operatorname{Rep}_{\alpha}(\Sigma,\operatorname{PSL}_{2}\operatorname{\mathbb{R}})$. If $\rho$ maps any non-peripheral simple closed curve on $\Sigma$ to a regular elliptic element, then $\rho$ is a DT representation. In particular, every totally elliptic representation is a DT representation.
The proof of Lemma 2.14 given in [Mar22] uses the characterization of DT representations in terms of Toledo number (given by (2.3)) and the additive properties of the Toledo number.
When $|\alpha|\in[2\pi,6\pi]$, the situation is not as simple as in Lemma 2.14, because there exist representations that map some non-peripheral simple closed curves on $\Sigma$ to elliptic elements, but which are not totally elliptic. An explicit example is provided by Palesi in [Pal14, Section 5.4]. We will nevertheless prove that if $|\alpha|\in[2\pi,6\pi]$ and $[\rho]\in\operatorname{Rep}_{\alpha}(\Sigma,\operatorname{PSL}_{2}\operatorname{\mathbb{R}})$ is the conjugacy class of a totally elliptic representation, then actually $|\alpha|\in\{2\pi,4\pi,6\pi\}$ and $\rho$ is orthogonal. We will use an argument of dynamical flavour that turns out to also cover the case of Lemma 2.14.
The pure mapping class group of $\Sigma$ acts on $\operatorname{Rep}_{\alpha}(\Sigma,\operatorname{PSL}_{2}\operatorname{\mathbb{R}})$ for any angle vector $\alpha\in(0,2\pi)^{4}$ (details can be found in [Mar22]). If $\rho$ is totally elliptic, then every point in the pure mapping class group orbit of $[\rho]$ is the conjugacy class of a totally elliptic representation. The orbit of $[\rho]$ is therefore bounded by Proposition 2.12. There are now two possibilities.
First, the orbit of $[\rho]$ may be finite. It turns out that finite orbits have been classified by Lisovyy–Tykhyy [LT14]. Their classification shows that any finite orbit in an $\alpha$-relative character variety $\operatorname{Rep}_{\alpha}(\Sigma,\operatorname{PSL}_{2}\operatorname{\mathbb{R}})$ is either an isolated point or belongs to a DT component. Recall from Corollary 2.11 that $\operatorname{Rep}_{\alpha}(\Sigma,\operatorname{PSL}_{2}\operatorname{\mathbb{R}})$ contains an isolated point if and only if $|\alpha|\in\{2\pi,4\pi,6\pi\}$, and the isolated point is always the conjugacy class of an orthogonal representation.
Secondly, we shall consider the case where the orbit of $[\rho]$ is an infinite bounded set. Infinite bounded orbits, in the context of 4-punctured spheres, have been studied by Cantat–Loray. A consequence of their work is the following statement. [^2] There is a small subtlety here. Cantat–Loray study the pure mapping class group action on the real points of the complex GIT quotient of $\operatorname{Hom}(\pi_{1}\Sigma,\mathrm{SL}_{2}\operatorname{\mathbb{C}})$ by $\mathrm{SL}_{2}\operatorname{\mathbb{C}}$ and not on the topological quotient of $\operatorname{Hom}(\pi_{1}\Sigma,\operatorname{PSL}_{2}\operatorname{\mathbb{R}})$ by $\operatorname{PSL}_{2}\operatorname{\mathbb{R}}$ like we do. However, since all the representations whose conjugacy class lies in an $\alpha$-relative character variety $\operatorname{Rep}_{\alpha}(\Sigma,\operatorname{PSL}_{2}\operatorname{\mathbb{R}})$ are reductive, Corollary 2.15 is indeed a consequence of Cantat–Loray’s results.
@@
If an $\alpha$-relative character variety $\operatorname{Rep}_{\alpha}(\Sigma,\operatorname{PSL}_{2}\operatorname{\mathbb{R}})$ contains an infinite bounded orbit, then $|\alpha|<2\pi$ or $|\alpha|>6\pi$ and the orbit is dense in the DT component of $\operatorname{Rep}_{\alpha}(\Sigma,\operatorname{PSL}_{2}\operatorname{\mathbb{R}})$.
Consequently, no matter if the orbit of $[\rho]$ is finite or infinite, $\rho$ is always either a DT or an orthogonal representation. This proves Theorem A for 4-punctured spheres.